OpenStudy (anonymous):

Solve the system using elimination. 5x = –25 + 5y 10y = 42 + 2x

OpenStudy (anonymous):

5x = -25 + 5y x = (-25 + 5y).2 same as dividing by 5, multiply both sides by .2 x = -5 + y Now, substitute your equation for x into the next equation. 10y = 42 + 2x 10y = 42 + 2(- 5 + y) 10y = 42 + (-10) + 2y put y's on one side, #'s on the other 10y -2y = 42 -10 8y = 32 y = 32/8 y = 4 Now go back to the equation you used to get x to find it's value. x = -5 + y x = -5 + 4 x = -1 Solution: (-1, 4) You can plug them into the equations and see if these work (they will)